# 3D Vision-Language Gaussian Splatting

### v1

@@ 1 Introduction
Recent methods (Engelmann et al., 2024; Qin et al., 2024; Shi et al., 2024; Zhou et al., 2024) in multi-modal 3D scene understanding have adopted the paradigm of embedding semantic representations into 3D representation for joint reconstruction training, and utilizing semantic knowledge distilled from off-the-shelf vision-language models to guide the training process. For example, OpenNeRF (Engelmann et al., 2024) integrates MipNeRF (Barron et al., 2021) method with OpenSeg model (Ghiasi et al., 2022), and LangSplat (Qin et al., 2024) employs both 3DGS (Kerbl et al., 2023) method and CLIP (Radford et al., 2021) model. These solutions rely on 2D supervision to learn a multi-modal (color and semantic) 3D scene representation, i.e., projecting the learned 3D representation back to 2D views for comparison with the input observations (input color images and corresponding 2D language maps inferred by aforementioned vision-language models).
@@ 2 Related Work
3D Neural Representations. Recent advancements in 3D scene representation have made significant progress, particularly with neural radiance fields (NeRF) (Mildenhall et al., 2020), which excel in novel view synthesis. However, NeRF’s reliance on a neural network for implicit scene representation can result in prolonged training and rendering times. Methods like Instant-NGP (Müller et al., 2022) speed up these processes through hash encoding, while approaches such as 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) use explicit neural representations for better alignment with GPU computations via differential tile rasterization. Additionally, MipNeRF (Barron et al., 2021) and Mip-Splatting (Yu et al., 2024) address aliasing issues, whereas DS-NeRF (Deng et al., 2022) and DG-Splatting (Chung et al., 2024) focus on sparse view reconstructions. In this paper, we utilize 3D Gaussian Splatting for 3D neural representations.
@@
Multi-modal 3D Scene Understanding. Existing methods for scene understanding learn multi-modal 3D representations from posed images, enabling rendering from novel viewpoints for tasks like open-vocabulary semantic segmentation. For NeRF-based approaches, LERF (Kerr et al., 2023) optimizes a language field alongside NeRF, using positions and physical scales to generate CLIP vectors. OpenNeRF (Engelmann et al., 2024) introduces a mechanism for obtaining novel camera poses, enhancing feature extraction. Among 3DGS-based methods, LangSplat (Qin et al., 2024) adopts a two-stage strategy for semantic feature acquisition, while GS-Grouping (Ye et al., 2024) extends Gaussian splatting for joint reconstruction and segmentation. Additionally, GOI (Qu et al., 2024) employs hyperplane division to select features for improved alignment.
@@ 3.1 Problem Statement
Gaussian splatting models have been expanded to incorporate language-embedding information densily describing the scene (Qin et al., 2024; Zhou et al., 2024; Qu et al., 2024), by adding a language feature vector $f^{i}\in\mathbb{R}^{d_{h}}$ (with $d_{h}$ feature size) to each 3D Gaussian and applying a similar rasterization process to this modality, i.e., the 2D semantic embedding $F$ at pixel $v$ can be expressed as:
@@
where $\mathcal{L}$ is the overall optimization objective. Besides, $F^{W}$ and $H^{W}$ are the predicted and ground-truth 2D language maps rendered using camera pose $W$ corresponding to image $I$ (for ease of readability, we drop the superscript $W$ in subsequent equations), and $\mathcal{L}_{\text{sem}}$ is a distance function for 2D semantic maps.
@@ 3.2 Semantic-aware Rasterization
Existing multi-modal rasterization methods (Qin et al., 2024; Zhou et al., 2024; Zuo et al., 2024) have largely drawn from color-based 3DGS (Kerbl et al., 2023). These rasterizers approach the rendering of 3D semantic information in a manner akin to RGB rendering (see Sec. 3.1), resulting in an insufficient focus on semantic-specific design. To address this gap, we propose a novel cross-modal rasterizer that emphasizes semantic-specific design, as illustrated in Fig. 2A and Fig. 2C.
@@
Another salient shortcoming is the simplistic adoption of $\alpha$-blending-based RGB splatting (Eq. 1) to rasterize language embeddings, only substituting the 3D color features $c$ with 3D semantic representations $f$ (Eq. 2). In other words, previous works use the opacity attribute of Gaussians to render not only RGB images but also language maps, c.f. Eq. 2 (either optimizing the opacities values over the two modalities (Shi et al., 2024; Zhou et al., 2024), or applying frozen opacity values from the RGB-only pre-training during semantic rasterization (Qin et al., 2024)). We argue that this approach fails to recognize that in phenomena involving semi-opaque media (e.g., glass, water) and complex light transport effects (e.g., direction-dependent scattering and glare), color opacity cannot be effectively translated to the semantic modality. In most cases, when Gaussians represent tangible elements of the scene (e.g., the surface or volume of an object), their semantic features should be splatted to 2D without any reduction in opacity. Conversely, intangible Gaussians (e.g., those simulating lighting effects like optical glares) should be excluded from semantic rasterization by enforcing their $\alpha$-blending weight close to zero.
@@ 3.3 Semantic-aware Camera View Blending
By combining the rasterization loss (Eq. 8) and the camera view blending loss (Eq. 11), the overall objective is represented as:
$\mathcal{L}_{\mathrm{overall}}=\mathbb{E}_{(I_{1},W_{1}),(I_{2},W_{2})\in T^{r}}[\mathcal{L}_{\mathrm{raster}}(W_{1})+\mathcal{L}_{\mathrm{raster}}(W_{2})+\lambda\mathcal{L}_{b}(W_{1},W_{2})],$ (12)
where $\lambda$ is a trade-off parameter.
@@ 4.1 Settings
Method Venue ramen figurines teatime kitchen avg.
@@
Method Venue ramen figurines teatime kitchen avg.
@@
Table 4: Ablation results of modality fusion on the LERF dataset, in terms of mIoU.
Method ramen figurines teatime kitchen avg.
No Fusion 57.8 69.6 54.2 51.5 58.3
@@
4.2 Results
Baselines. For a fair comparison, we select the latest works on open-vocabulary 3D scene understanding: Feature-3DGS (Zhou et al., 2024), LEGaussians (Shi et al., 2024), LangSplat (Qin et al., 2024), GS-Grouping (Ye et al., 2024), GOI (Qu et al., 2024), and FMGS (Zuo et al., 2024), all of which are based on 3DGS (Kerbl et al., 2023), and employ the same segmentation and vision-language models mentioned in Sec. 4.1 to extract ground truth language features. It is important to note that FMGS (Zuo et al., 2024) does not report mIoU results on the LERF dataset and is also not open-sourced, so its mIoU results are not listed.
Qualitative Evaluation. To facilitate qualitative comparisons, we contrast our method with several baseline approaches, including LEGaussian (Shi et al., 2024), LangSplat (Qin et al., 2024), and GOI (Qu et al., 2024). Fig. 4 and Fig. 5 illustrate the semantic segmentation results on the LERF dataset, while Figure 6 presents the localization results on the same dataset. For the localization task, the black bounding boxes with dotted lines represent the ground truth ranges, while the red dots indicate predictions from various methods. Besides, segmentation comparisons on the 3D-OVS dataset are displayed in Fig. 7 and Fig. 8. It is clear that our proposed method outperforms the other approaches and is closer to the ground-truth, particularly in reflective and translucent areas. Additional results are displayed in the appendix.
4.3 Ablation Studies
Ablation on cross-modal rasterizer. Tab. 4 presents an ablation study on the data-fusion of the rasterizer. In addition to no fusion like LEGaussians (Shi et al., 2024), we also incorporate single-layer MLP and cross-attention for modality Fusion to enhance semantic features for comparison. The results indicate that self-attention modality fusion is the most effective method. Besides, single-layer MLP and cross-attention modality fusion do not demonstrate any advantage over the no fusion condition. In Tab. 5, we highlight the benefits of our proposed smoothed semantic indicator. We compare it to previous methods for rasterizing semantic embeddings, specifically by applying the color opacity $o^{i}$ to the language modality (Qin et al., 2024; Zhou et al., 2024; Zuo et al., 2024). The results indicate that using a separate attribute for rasterizing language embeddings is clearly advantageous. However, this new attribute cannot be naively fixed, e.g., to 1 or 0.5 for all Gaussians. Its values should be learned by the model through 2D supervision, which is why we introduce a learnable indicator for the $\alpha$-blending of language modality.
Table 5: Ablation results of smoothed semantic indicator on the LERF dataset, in terms of mIoU.
Method ramen figurines teatime kitchen avg.
Fixed to 0.5 52.2 70.9 50.6 50.8 56.1
@@ 4.3 Ablation Studies
Rotation Translation SSIM ramen figurines teatime kitchen avg.
@@
Table 7: Ablation results of interpolation ratio sampling on the LERF dataset, in terms of mIoU.
Method ramen figurines teatime kitchen avg.
Fixed to 0.5 58.8 72.6 56.4 52.5 60.1
Fixed to 0.75 56.3 71.0 50.5 51.7 57.4
Fixed to 1.0 55.2 68.3 48.5 51.0 55.8
Even Distribution 58.3 71.1 55.4 51.2 59.0
Gaussian Distribution 57.5 71.3 53.9 50.7 58.4
Beta Distribution (Ours) 61.4 73.5 58.1 54.8 62.0
In Tab. 7, we present an ablation study on the interpolation ratio $k$. In addition to fixed values of $0.5$, $0.75$, and $1.0$, we also include experiments based on Even Distribution (U$(0,1)$), Gaussian Distribution (N$(0,1)$), and our Beta Distribution (Beta(0.2,0.2)0.20.2(0.2,0.2)( 0.2 , 0.2 )). The results demonstrate that a balanced ratio is crucial for acquiring better representations, and that slight disturbances can be beneficial as well.
@@ A.3 Additional Ablation Studies
Figure 9: Qualitative semantic segmentation comparisons on the bonsai scene of the Mip-NeRF 360 dataset.
Figure 10: Qualitative semantic segmentation comparisons on the room scene of the Mip-NeRF 360 dataset.
@@
Ablation of Semantic Indicator.
@@
Ablation of Camera View Blending.
We present an ablation study of various camera view blending strategies in Tab. 13 using the 3D-OVS dataset, which includes three variants: Rotation, Translation, and SSIM. Our proposed method is represented in the last row of the table. The results indicate that both Slerp-based rotation blending and Lerp-based translation blending enhance performance. Furthermore, the regularization provided by SSIM is crucial for managing the extent of blending.
@@
Table 15: Ablation results of the parameters of Beta distribution on the LERF dataset, in terms of mIoU.
Method ramen figurines teatime kitchen avg.
@@
In Tab. 14, we further present an ablation study on the interpolation ratio used within the camera view blending scheme, performed on the 3D-OVS dataset. Alongside fixed values of $0.5$, $0.75$, and $1.0$, we also include experiments utilizing Even Distribution (U$(0,1)$), Gaussian Distribution (N$(0,1)$), and our Beta Distribution (Beta(0.2,0.2)0.20.2(0.2,0.2)( 0.2 , 0.2 )). The results indicate that a balanced ratio is essential for achieving better representations, and that minor disturbances can also yield beneficial effects.

### v2

@@ 1 Introduction
Recent methods (Engelmann et al., 2024; Qin et al., 2024; Shi et al., 2024; Zhou et al., 2024b; Yu et al., 2024a) in multi-modal 3D scene understanding have adopted the paradigm of embedding semantic representations into 3D representation for joint reconstruction training, and utilizing semantic knowledge distilled from off-the-shelf vision-language models to guide the training process. For example, OpenNeRF (Engelmann et al., 2024) integrates MipNeRF (Barron et al., 2021) method with OpenSeg model (Ghiasi et al., 2022), and LangSplat (Qin et al., 2024) employs both 3DGS (Kerbl et al., 2023) method and CLIP (Radford et al., 2021) model. These solutions rely on 2D supervision to learn a multi-modal (color and semantic) 3D scene representation, i.e., projecting the learned 3D representation back to 2D views for comparison with the input observations (input color images and corresponding 2D language maps inferred by aforementioned vision-language models).
@@ 2 Related Work
3D Neural Representations. Recent advancements in 3D scene representation have made significant progress, particularly with neural radiance fields (NeRF) (Mildenhall et al., 2020), which excel in novel view synthesis. However, NeRF’s reliance on a neural network for implicit scene representation can result in prolonged training and rendering times. Methods like Instant-NGP (Müller et al., 2022) speed up these processes through hash encoding, while approaches such as 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) use explicit neural representations for better alignment with GPU computations via differential tile rasterization. Additionally, MipNeRF (Barron et al., 2021) and Mip-Splatting (Yu et al., 2024b) address aliasing issues, whereas DS-NeRF (Deng et al., 2022) and DG-Splatting (Chung et al., 2024) focus on sparse view reconstructions. In this paper, we utilize 3D Gaussian Splatting for 3D neural representations.
@@
Multi-modal 3D Scene Understanding. Existing methods for scene understanding learn multi-modal 3D representations from posed images, enabling rendering from novel viewpoints for tasks like open-vocabulary semantic segmentation. For NeRF-based approaches, LERF (Kerr et al., 2023) optimizes a language field alongside NeRF, using positions and physical scales to generate CLIP vectors. OpenNeRF (Engelmann et al., 2024) introduces a mechanism for obtaining novel camera poses, enhancing feature extraction. Among 3DGS-based methods, LangSplat (Qin et al., 2024) adopts a two-stage strategy for semantic feature acquisition, while GS-Grouping (Ye et al., 2024) extends Gaussian splatting for joint reconstruction and segmentation. Additionally, GOI (Qu et al., 2024) employs hyperplane division to select features for improved alignment. Moreover, HUGS Zhou et al. (2024a) enables holistic 3D scene understanding by integrating 2D semantics and flow with 3D tracking, effectively lifting them into the 3D space.
@@ 3.1 Problem Statement
Gaussian splatting models have been expanded to incorporate language-embedding information densely describing the scene (Qin et al., 2024; Zhou et al., 2024b; Qu et al., 2024) by adding a language feature vector $f^{i}\in\mathbb{R}^{d_{h}}$ (with $d_{h}$ feature size) to each 3D Gaussian and similarly rasterizing this modality. I.e., a 2D semantic embedding $F$ at pixel $v$ can be expressed as:
@@
where $\mathcal{L}$ is the overall optimization objective. Besides, $F^{W}$ and $H^{W}$ are the predicted and ground-truth 2D language maps rendered using camera pose $W$ corresponding to image $I$ (for ease of readability, we drop the superscript $W$ in subsequent equations), and $\mathcal{L}_{\text{sem}}$ is a distance function for 2D semantic maps (e.g., L1 distance).
@@ 3.2 Semantic-aware Rasterization
Existing multi-modal rasterization methods (Qin et al., 2024; Zhou et al., 2024b; Zuo et al., 2024) have largely drawn from color-based 3DGS (Kerbl et al., 2023). These rasterizers approach the rendering of 3D semantic information in a manner akin to RGB rendering (see Sec. 3.1), resulting in an insufficient focus on semantic-specific design. To address this gap, we propose a novel cross-modal rasterizer that emphasizes semantic-specific design, as illustrated in Fig. 2A and Fig. 2C.
@@
Another salient shortcoming is the simplistic adoption of $\alpha$-blending-based RGB splatting (Eq. 1) to rasterize language embeddings, only substituting the 3D color features $c$ with 3D semantic representations $f$ (Eq. 2). In other words, previous works use the opacity attribute of Gaussians to render not only RGB images but also language maps, c.f. Eq. 2 (either optimizing the opacities values over the two modalities (Shi et al., 2024; Zhou et al., 2024b), or applying frozen opacity values from the RGB-only pre-training during semantic rasterization (Qin et al., 2024)). We argue that this approach fails to recognize that in phenomena involving semi-opaque media (e.g., glass, water) and complex light transport effects (e.g., direction-dependent scattering and glare), color opacity cannot be effectively translated to the semantic modality. In most cases, when Gaussians represent tangible elements of the scene (e.g., the surface or volume of an object), their semantic features should be splatted to 2D without any reduction in opacity. Conversely, intangible Gaussians (e.g., those simulating lighting effects like optical glares) should be excluded from semantic rasterization by enforcing their $\alpha$-blending weight close to zero.
@@ 3.3 Semantic-aware Camera View Blending
By combining the rasterization loss (Eq. 8) and the camera view blending loss (Eq. 11) according to a trade-off hyper-parameter $\lambda$, the overall objective is represented as:
$\mathcal{L}_{\mathrm{overall}}=\mathbb{E}_{(I_{1},W_{1}),(I_{2},W_{2})\in T^{r}}[\mathcal{L}_{\mathrm{raster}}(W_{1})+\mathcal{L}_{\mathrm{raster}}(W_{2})+\lambda\mathcal{L}_{b}(W_{1},W_{2})].$ (12)
@@ 4.1 Settings
4.2 Results
Baselines. For a fair comparison, we select the latest works on open-vocabulary 3D scene understanding: Feature-3DGS (Zhou et al., 2024b), LEGaussians (Shi et al., 2024), LangSplat (Qin et al., 2024), GS-Grouping (Ye et al., 2024), GOI (Qu et al., 2024), and FMGS (Zuo et al., 2024), all of which are based on 3DGS (Kerbl et al., 2023), and employ the same segmentation and vision-language models mentioned in Sec. 4.1 to extract ground truth language features. It is important to note that FMGS (Zuo et al., 2024) does not report mIoU results on the LERF dataset and is also not open-sourced, so its mIoU results are not listed.
Qualitative Evaluation. To facilitate qualitative comparisons, we contrast our method with several baseline approaches, including LEGaussian (Shi et al., 2024), LangSplat (Qin et al., 2024), and GOI (Qu et al., 2024). Fig. 4 and Fig. 5 illustrate semantic segmentation results on LERF data, while Fig. 6 presents the localization results on the same dataset. For the localization task, the black bounding boxes with dotted lines represent the ground truth ranges, while the red dots indicate predictions from various methods. Besides, segmentation comparisons on the 3D-OVS dataset are displayed in Fig. 7 and Fig. 8. It is clear that our proposed method outperforms the other approaches and is closer to the ground-truth, particularly in reflective and translucent areas. Additional results are displayed in the appendix.
4.3 Ablation Studies
Ablation on cross-modal rasterizer. Tab. 4.A presents an ablation study on the data-fusion of the rasterizer. In addition to no fusion like LEGaussians (Shi et al., 2024), we also incorporate single-layer MLP and cross-attention for modality Fusion to enhance semantic features for comparison. The results indicate that self-attention modality fusion is the most effective method. Besides, single-layer MLP and cross-attention modality fusion do not demonstrate any advantage over the no fusion condition. In Tab. 4.B, we highlight the benefits of our proposed smoothed semantic indicator. We compare it to previous methods for rasterizing semantic embeddings, specifically by applying the color opacity $o^{i}$ to the language modality (Qin et al., 2024; Zhou et al., 2024b; Zuo et al., 2024). The results indicate that using a separate attribute for rasterizing language embeddings is clearly advantageous. However, this new attribute cannot be naively fixed, e.g., to 1 or 0.5 for all Gaussians. Its values should be learned by the model through 2D supervision, which is why we introduce a learnable indicator for the $\alpha$-blending of language modality.
In Tab. 4.C, we present an ablation study on the interpolation ratio $k$. In addition to fixed values of $0.5$, $0.75$, and $1.0$, we also include experiments based on Even Distribution (U$(0,1)$), Gaussian Distribution (N$(0,1)$), and our Beta Distribution (Beta(0.2,0.2)0.20.2(0.2,0.2)( 0.2 , 0.2 )). The results demonstrate that a balanced ratio is crucial for acquiring better representations, and that slight disturbances can be beneficial as well.
@@ 4.3 Ablation Studies
Method Venue ramen teatime figurines kitchen avg.
@@
Method Venue ramen teatime figurines kitchen avg.
@@
Table 4: Ablation results of various contributions on the LERF dataset, in terms of mIoU.
Method ramen teatime figurines kitchen avg.
(A) Fusion module No Fusion 57.8 69.6 54.2 51.5 58.3
@@
(B) Indicator value Fixed to 0.5 52.2 70.9 50.6 50.8 56.1
@@
(C) Interpolation ratio sampling Fixed to 0.5 58.8 72.6 56.4 52.5 60.1
Fixed to 0.75 56.3 71.0 50.5 51.7 57.4
Fixed to 1.0 55.2 68.3 48.5 51.0 55.8
Even Distribution 58.3 71.1 55.4 51.2 59.0
Gaussian Distribution 57.5 71.3 53.9 50.7 58.4
Beta Distribution (Ours) 61.4 73.5 58.1 54.8 62.0
@@
Rotation Translation SSIM ramen teatime figurines kitchen avg.
@@ 5 Conclusion
Ethical Statement
We have carefully read the ICLR Code of Ethics. We are sure that all studies and procedures described in the paper follow the ethical rules in the academia. Our research does not involve human subjects, and also does not contain any potentially harmful insights, methodologies and applications, Furthermore, we do not expect any direct discrimination/bias/fairness concerns, or privacy/security issues relating to this paper.
Reproducibility Statement
All key technical details and quantitative results are provided in the main paper and Appendix, enabling researchers and practitioners in this field to reproduce our work. Additionally, all datasets used in this research are publicly available to the community. We will also release the source code of our method upon acceptance of the paper.
@@ A.1 Further Implementation Details
For a larger version of Sub-figs. 2.C-D representing our fusion module and augmentation module respectively, we refer the readers to Fig. 9.
In the ablation on cross-modal rasterizer (Sec. 4.3), the cross-attention fusion tested in Tab. 4.A is performed by using $f^{i}$ as query and $c^{i}$ as both key and value, whereas the adopted self-attention performs cross-modality fusion by using $c^{i}\oplus f^{i}$ as query, key, and value.
Figure 9: Larger version of Sub-figs. 2.C-D describing our fusion and augmentation modules.
@@ A.2 Evaluation on Additional Mip-NeRF 360 Dataset
A.2.2 Qualitative Results.
Figure 10: Qualitative semantic segmentation comparisons on the bonsai scene of Mip-NeRF 360.
Figure 11: Qualitative semantic segmentation comparisons on the room scene of Mip-NeRF 360.
Table 9: Ablation results of three key contributions on the LERF dataset, in terms of mIoU.
modal. fus. sem. indic. view blend. ramen teatime figurines kitchen avg.
$\times$ $\times$ $\times$ 53.3 66.9 46.3 45.3 53.0
$\times$ $\times$ $\checkmark$ 54.7 68.5 47.7 45.8 54.2
$\times$ $\checkmark$ $\times$ 55.4 67.2 47.8 48.9 54.8
$\checkmark$ $\times$ $\times$ 54.8 67.5 48.2 48.6 54.8
$\times$ $\checkmark$ $\checkmark$ 57.8 69.6 54.2 51.5 58.3
$\checkmark$ $\times$ $\checkmark$ 55.9 71.0 54.2 51.3 58.1
$\checkmark$ $\checkmark$ $\times$ 55.4 68.2 48.8 50.9 55.8
$\checkmark$ $\checkmark$ $\checkmark$ 61.4 73.5 58.1 54.8 62.0
Table 10: Ablation results of three key contributions on the 3D-OVS dataset, in terms of mIoU.
modal. fus. sem. indic. view blend. bed bench room sofa lawn avg.
$\times$ $\times$ $\times$ 92.7 94.3 94.1 90.5 96.0 93.5
$\times$ $\times$ $\checkmark$ 93.0 94.6 94.9 91.6 96.1 94.0
$\times$ $\checkmark$ $\times$ 93.6 95.1 95.0 92.2 96.0 94.4
$\checkmark$ $\times$ $\times$ 93.9 95.2 95.3 91.8 96.2 94.5
$\times$ $\checkmark$ $\checkmark$ 94.5 95.6 95.3 93.0 96.5 95.0
$\checkmark$ $\times$ $\checkmark$ 94.7 95.2 95.0 92.7 96.4 94.8
$\checkmark$ $\checkmark$ $\times$ 92.9 94.6 95.2 91.5 96.1 94.1
$\checkmark$ $\checkmark$ $\checkmark$ 96.8 97.3 97.7 95.5 97.9 97.1
@@
Table 13: Comparison between data-augmentation strategies, over downstream open-vocabulary semantic segmentation (mIoU) on LERF scenes. Note that the original training sets contain between 124 and 297 images per scene. For the offline augmentation using 3DGS-rendered images and off-the-shelf open-vocabulary models (SAM/CLIP), 120 novel views are added, with their viewpoints sampled via interpolation.
Method ramen teatime figurines kitchen avg.
w/o data augmentation 53.3 66.9 46.3 45.3 53.0
w/ data from 3DGS & SAM/CLIP 55.2 68.1 49.5 48.2 55.3
w/ view blending 61.4 73.5 58.1 54.8 62.0
@@
Table 16: Ablation results of the parameters of Beta distribution on LERF data, in terms of mIoU.
Method ramen teatime figurines kitchen avg.
@@
A.3 Additional Ablation Studies
A.3.1 High-level Ablation of Proposed Contributions.
We first provide some high-level ablation studies Tab. 10 and Tab. 10, highlighting the impact of each of our key contributions, i.e., our proposed modality fusion (“ modal. fus.”), semantic indicator (“ sem. indic.”), and camera-view blending (“view blend.”) . The results demonstrate that all three components are crucial in improving the overall performance.
A.3.2 Ablation of Modality Fusion
A.3.3 Ablation of Semantic Indicator
A.3.4 Ablation of Camera View Blending
We first provide further justifications for our view-blending augmentation technique. Instead of our mix-up strategy, one could apply the selected off-the-shelf language-feature extractor (e.g., CLIP) to generate new pseudo-ground-truth data. This could be done either:
  • online, i.e., using the model’s predicted color image as input to the language-feature extractor in order to generate the target language map to evaluate the one rendered by the model.
  • offline, i.e., using a pre-trained color-only 3DGS model of the scene to generate novel views, pass these views to the feature extractor, and use all the resulting data to train the language-embedded representation.
However, the computational cost of running the selected language-feature extractor (CLIP) is too heavy for online usage, and the offline strategy also comes with several downsides.
First, one of our contributions is the joint CUDA-based rasterization function, which enables the fast rendering of both modalities (color and language). While most prior works Qin et al. (2024) train their model in two phases (a first phase to train a color-only 3DGS model, then a second phase to extend it with language information), we leverage our custom rasterizer to efficiently optimize our vision-language model in a single training phase (see Table 8 highlighting our training and rendering efficiency). The downside of this faster training is that, unlike prior solutions, we do not have access to a pre-trained 3DGS model of the target scene. Adding a pre-training phase for the color-only model would be a significant computational overhead. Second, the aforementioned strategy (i.e., rendering additional views using a pre-trained 3DGS) only adds a fixed/limited number of images to the training set (c.f. offline rendering before the training starts), whereas our augmentation scheme provides new randomized language maps at every training iteration (c.f. online augmentation).
We quantitatively compared the suggested offline augmentation strategy to ours in a new experiment shared in Tab. 16, which confirms that our augmentation technique does indeed result in a more robust representation w.r.t. downstream semantic tasks. Note that in future work, it would be interesting to investigate how both strategies could be optimally interleaved (e.g., by pretraining our model with mix-up augmentation; then, after some iterations, augmenting the training data with 3DGS/CLIP-rendered pairs).
Additionally, we present an ablation study of various camera view blending strategies in Tab. 16 using the 3D-OVS dataset, which includes three variants: Rotation, Translation, and SSIM. Our proposed method is represented in the last row of the table. The results indicate that both Slerp-based rotation blending and Lerp-based translation blending enhance performance. Furthermore, the regularization provided by SSIM is crucial for managing the extent of blending.
In Tab. 16, we further present an ablation study on the interpolation ratio used within the camera view blending scheme, performed on 3D-OVS data. Alongside fixed values of $0.5$, $0.75$, and $1.0$, we also include experiments utilizing Even Distribution (U$(0,1)$), Gaussian Distribution (N$(0,1)$), and our Beta Distribution (Beta(0.2,0.2)0.20.2(0.2,0.2)( 0.2 , 0.2 )). The results indicate that a balanced ratio is essential for achieving better representations, and that minor disturbances can also yield beneficial effects.